SLR has set a target of 20% annual rate of return on investments. The fixed assets employed on one of their products, the X2, amount to £21,000,000. Working capital investment for the X2 is £4,000,000. Total operating costs to be incurred on the X2 for next year are forecast to be £8,000,000. SLR uses cost plus pricing.

 

What is the required percentage mark up on the costs for X2?


1
Expert's answer
2021-12-02T10:40:28-0500

Solution:

The required percentage markup on the costs for X2:

Mark-up = "\frac{Selling\; price - Cost }{Selling\; price} \times 100\%"


Selling price = Cost plus 20%

Total costs = 8,000,000 + 21,000,000 = 29,000,000

Selling price = 120"\%" "\times" 29,000,000 = 34,800,000

Mark-up = "\frac{34,800,000 - 29,000,000 }{34,800,000} \times 100\% =" 16.67%


The required percentage markup on the costs for X2 = 16.67%